OpenStudy (anonymous):

solve [1-(x/x+1)]/[2-(x-1/x)]

OpenStudy (tkhunny):

There is nothing to solve. Maybe "simplify"? Is this what you mean? \(\dfrac{1-\dfrac{x}{x+1}}{2-\dfrac{x-1}{x}}\) It is NOT what you have provided.

OpenStudy (anonymous):

what does "solve" mean in this context?

OpenStudy (dan815):